# On a discrete version of the position-momentum commutation relation

### v1

@@ I Introduction
In quantum mechanics, a pair AA and $B$ of Hermitian operators on a Hilbert space is canonical if they satify the relation $A,B]\!=\!\mathrm{i}\hbar$, where [A,B]=ABBA[A,B]\!=\!AB-\!BA, $h\!=\!2\pi\hbar$ is the Plank constant, and $\mathrm{i}\hbar\equiv\mathrm{i}\hbar I$ is the identity operator multiplied by $\mathrm{i}\hbar$. In the finite-dimensional case, such a pair does not exist because the trace of the two sides of $A,B]\!=\!\mathrm{i}\hbar$ is different. In [6], the authors show that certain canonical pairs can be obtained by relaxing the definition, namely by imposing the relation $A,B]\!=\!\mathrm{i}\hbar$ to be satisfied only on a proper subspace of the Hilbert space. Some references concerning possible applications of this new definition are presented in [6].
@@ V Discrete version of the position-momentum commutation relation
$\begin{array}[]{l}\langle n|\hat{\mathfrak{C}}|m\rangle\!=\!\frac{1}{d}\sum\limits_{k=-s}^{s}k(n\!-\!m)\,\mathrm{Exp}[-\frac{2\pi\mathrm{i}}{d}k(n-m)]-{\rm i}\mbox{\small$\frac{d}{2\pi}$}\delta_{nm}\end{array}$ (27)
@@
Let $\lambda_{0}$, $\lambda_{1}$, …, $\lambda_{d-1}$ be the eigenvalues of $\hat{\mathfrak{C}}\!=\![\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}$, considered in the increasing order of their modulus ($|\lambda_{1}|\leq|\lambda_{2}|\leq...\leq|\lambda_{d}|$), and let $\varphi_{0}$, $\varphi_{1}$, … , $\varphi_{d-1}$ be the corresponding eigenfunctions, that is
@@ XIII Mehta states
It is known that, in the continuous-variable case, the eigenstates of the quantum harmonic oscillator $\hat{H}\!=\!\frac{{p}^{2}\!+\!\hat{q}^{2}}{2}$ can be defined in terms of Hermite polynomials as $\Psi_{n}\!:\!(-\infty,\infty)\!\rightarrow\mathbb{R}$,
@@ XIV Harper states
$\hat{H}\!=\!\frac{{p}^{2}\!+\!\hat{q}^{2}}{2}\!=\!\frac{{p}^{2}\!+\!\hat{F}\hat{p}^{2}\hat{F}^{\dagger}}{2}.$ (80)

### v2

@@ I Introduction
In quantum mechanics, a pair $A$ and $B$ of Hermitian operators on a Hilbert space is canonical if they satify the relation $A,B]\!=\!\mathrm{i}\hbar$, where [A,B]=ABBA[A,B]\!=\!AB-\!BA, $h\!=\!2\pi\hbar$ is the Plank constant, and $\mathrm{i}\hbar\equiv\mathrm{i}\hbar I$ is the identity operator multiplied by $\mathrm{i}\hbar$. In the finite-dimensional case, such a pair does not exist because the trace of the two sides of $A,B]\!=\!\mathrm{i}\hbar$ is different. In [6], the authors show that certain canonical pairs can be obtained by relaxing the definition, namely by imposing the relation $A,B]\!=\!\mathrm{i}\hbar$ to be satisfied only on a proper subspace of the Hilbert space. Some references concerning possible applications of this new definition are presented in [6].
@@ V Discrete version of the position-momentum commutation relation
$\begin{array}[]{l}\langle n|\hat{\mathfrak{C}}|m\rangle\!=\!\frac{1}{d}\sum\limits_{k=-s}^{s}k(n\!-\!m)\,\mathrm{Exp}[\frac{2\pi\mathrm{i}}{d}k(n-m)]-{\rm i}\mbox{\small$\frac{d}{2\pi}$}\delta_{nm}\end{array}$ (27)
@@
Let $\lambda_{0}$, $\lambda_{1}$, …, $\lambda_{d-1}$ be the eigenvalues of $\hat{\mathfrak{C}}\!=\![\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}$, considered in the increasing order of their modulus ($|\lambda_{0}|\leq|\lambda_{1}|\leq...\leq|\lambda_{d-1}|$), and let $\varphi_{0}$, $\varphi_{1}$, … , $\varphi_{d-1}$ be the corresponding eigenfunctions, that is
@@ XIII Mehta states
It is known that, in the continuous-variable case, the eigenstates of the quantum harmonic oscillator $\hat{H}\!=\!\frac{\hat{p}^{2}\!+\!\hat{q}^{2}}{2}$ can be defined in terms of Hermite polynomials as $\Psi_{n}\!:\!(-\infty,\infty)\!\rightarrow\mathbb{R}$,
@@ XIV Harper states
$\hat{H}\!=\!\frac{\hat{p}^{2}\!+\!\hat{q}^{2}}{2}\!=\!\frac{\hat{p}^{2}\!+\!\hat{F}\hat{p}^{2}\hat{F}^{\dagger}}{2}.$ (80)